OpenStudy (anonymous):

What is the average value of f(x)=sqrt 64-x^2 over the interval 0<=x<=8

OpenStudy (anonymous):

how should i approach this problem?

OpenStudy (anonymous):

the way you do this is you plug in 0 into the function then you plug in 8 and you take the average so we have sqrt(64-0)=8 and sqrt(64-8) = sqrt(56) so then 8+sqrt(56)/8 im pretty sure

OpenStudy (anonymous):

which would be square root of 56 whatever that number is..

OpenStudy (anonymous):

that doesnt sound right...you can't just cancel the 8

OpenStudy (anonymous):

it isnt cancelling it... the 8 is due to the interval that just worked out that way because they chose the number 8 and the number 64

OpenStudy (anonymous):

so the answer is \[\sqrt{56}\]

OpenStudy (anonymous):

The average value of a function is given by: \[1/(b-a) \int\limits_{a}^{b}f(x) dx\] In your problem, a=0, b=8.

OpenStudy (anonymous):

yeah but that doesnt really work..i already tried it...can you walk me through it

OpenStudy (anonymous):

nvrmind...im just gonna use Reiman Sum